In a study of the mating calls of the tree toad Hyla, Littlejohn (1965) found the duration
of the call in a sample of 29 observations to have a mean of 189 ms (milliseconds) and a
standard deviation of 31.0 ms.
(a) Construct the 95 % confidence interval to the mean call duration (5)
(b) Construct the 90 % confidence interval to the mean call duration (5)
(c) Interpret what these confidence intervals signify and explain why they are different
1
Expert's answer
2020-04-20T11:15:34-0400

a) 95% confidence interval

CI =mean ± margin error (E)

E="Z_c\frac {\sigma} {\sqrt n}"

="1.96*\frac {31}{\sqrt {29}}"

=11.2828

CI=189±11.2828

=(177.7172,200.2828)

b) 90% confidence interval

="1.64*\frac {31}{\sqrt {29}}"

=9.4408

CI= 189±9.4408

=(179.5592, 198.4408)

c) These confidence intervals signify the range in which the population parameter lies. The values are different because the confidence levels are different. A higher confidence level has a wider confidence width.